Ahmer walks 795m east and 55m north 80m east find the displacement and distance?
1
Expert's answer
2020-03-06T10:26:18-0500

The distance


"d=795+55+80=930\:\rm m."

The displacement


"s=\sqrt{(795+80)^2+55^2}=877\:\rm m."